OpenStudy (anonymous):

what values for Ø(0 ≤Ø ≤ 2 pi ) satisfy the equation 3 sin Ø= sin Ø-1

OpenStudy (anonymous):

@goformit100 @geekfromthefutur @ofmice&men

OpenStudy (anonymous):

@foreverjennnicole @no;biuvjh,mgj

OpenStudy (kagıtucak):

3sina=sina-1 2sina=-1 sina=-1/2 sinüs is minus in third and fourth quadrant and sin60=1/2 so 60+180=240 and 360-60=300 degrees are the answers.

OpenStudy (anonymous):

the possible answers are a: pi/3 , 5pi/3 b: 2pi/3 , 4pi/3 c: pi/6 , 3pi/6 d: 7pi/6 , 11pi/6

OpenStudy (kagıtucak):

I made a terrible mistake sin60=3^(1/2)/2 sin30=1/2 :( Thus 180+30=210 and 360-30=330 are the answers so it is d :)

OpenStudy (anonymous):

thank you QAQ would you mind sticking around and helping me with the rest?

OpenStudy (kagıtucak):

ı can see them

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

the next one is the same but with a different equation. tan^2 Ø = -3/2 sec Ø and the possible answers are a: pi/2 , 5pi/2 b: 3pi/4 , 5pi/4 c: pi/3 , 5pi/3 d: 2pi/3 , 4pi/3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

on this page http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf you'll see that \[\Large \tan^2(\theta) + 1 = \sec^2(\theta)\] (see page 2)

jimthompson5910 (jim_thompson5910):

solve for tan^2 to get \[\Large \tan^2(\theta) = \sec^2(\theta) - 1\]

jimthompson5910 (jim_thompson5910):

Then plug that into the equation to get \[\Large \tan^2(\theta) = -\frac{3}{2}\sec(\theta)\] \[\Large \sec^2(\theta) - 1 = -\frac{3}{2}\sec(\theta)\]

jimthompson5910 (jim_thompson5910):

do you know how to solve from here?

OpenStudy (anonymous):

no Q_Q i dont know how to do any of this thats why i'm here

jimthompson5910 (jim_thompson5910):

ok let z = sec(theta) that means z^2 = sec^2(theta)

jimthompson5910 (jim_thompson5910):

and that allows us to replace every copy of sec(theta) with z to get z^2 - 1 = (-3/2)*z

jimthompson5910 (jim_thompson5910):

are you able to solve z^2 - 1 = (-3/2)*z for z?

OpenStudy (anonymous):

no... sorry...

jimthompson5910 (jim_thompson5910):

z^2 - 1 = (-3/2)*z 2z^2 - 2 = -3z ... multiply everything by 2 to clear out the fraction 2z^2 - 2 + 3z = 0 2z^2 + 3z - 2 = 0 Now use the quadratic formula to solve for z

OpenStudy (anonymous):

how would this get any of the possible answers? the possible answers are all fractions with pi

jimthompson5910 (jim_thompson5910):

we'll get there, don't worry

jimthompson5910 (jim_thompson5910):

what solutions do you get in terms of z?

OpenStudy (anonymous):

i dont know the quadratic formula...

jimthompson5910 (jim_thompson5910):

the formula (and examples using it) is given here http://www.purplemath.com/modules/quadform.htm

jimthompson5910 (jim_thompson5910):

in this case, a = 2, b = 3, c = -2

OpenStudy (anonymous):

1.5/4 and 7.5/4

jimthompson5910 (jim_thompson5910):

You should get \[\Large z=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large z=\frac{-3 \pm \sqrt{(3)^2-4(2)(-2)}}{2(2)}\] \[\Large z=\frac{-3 \pm \sqrt{25}}{4}\] \[\Large z=\frac{-3 \pm 5}{4}\] \[\Large z=\frac{-3 + 5}{4} \ \text{ or } \ z=\frac{-3 - 5}{4}\] \[\Large z=\frac{2}{4} \ \text{ or } \ z=\frac{-8}{4}\] \[\Large z=\frac{1}{2} \ \text{ or } \ z=-2\]

jimthompson5910 (jim_thompson5910):

if z = 1/2, and z = sec(theta), then z = 1/2 sec(theta) = 1/2 theta = ????

OpenStudy (anonymous):

i dont know Q_Q

jimthompson5910 (jim_thompson5910):

use the identity sheet I gave you secant is what in terms of cosine?

OpenStudy (anonymous):

i looked but i cant find it...

jimthompson5910 (jim_thompson5910):

look under reciprocal identities

jimthompson5910 (jim_thompson5910):

near the top of page 2

OpenStudy (anonymous):

sec=1/cos???

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so if sec(theta) = 1/2, then taking the reciprocal of both sides gives 1/sec(theta) = 2/1 cos(theta) = 2 is it possible for cos(theta) = 2 ?

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

actually no it's not possible because the range of cosine is -1 to 1 1 is the largest value cosine can be

jimthompson5910 (jim_thompson5910):

cos(theta) = 2 has no solutions

jimthompson5910 (jim_thompson5910):

if z = -2, then... z = -2 sec(theta) = -2 1/cos(theta) = -2 cos(theta) = -1/2 use a unit circle to find the value of theta that makes that equation true

OpenStudy (anonymous):

i still dont understand any of this

jimthompson5910 (jim_thompson5910):

so none of this makes any sense?

OpenStudy (anonymous):

i'm trying to understand but...

jimthompson5910 (jim_thompson5910):

if so, then perhaps you're in the wrong lesson?

OpenStudy (anonymous):

this class is mandatory... its required to graduate.. and they went and stuffed difficult trig into an algebra class

OpenStudy (anonymous):

the class is already so difficult that i am FAILING. i understand none of it and my teacher teaches all of the other math classes for the school so shes literally useless

jimthompson5910 (jim_thompson5910):

completely unfair since trig and algebra are somewhat different classes

jimthompson5910 (jim_thompson5910):

there's no way to transfer?

OpenStudy (anonymous):

nope. not that my mom would let me anyways. she doesnt care. all she does is yell at me and tell me to "do better"

jimthompson5910 (jim_thompson5910):

so since this is an algebra class (I thought it was a trig class), I'm guessing this is the first time you've seen the quadratic formula?

OpenStudy (anonymous):

they had the quadratic formula in algebra 1 but i havent used it since so i had forgotten what it was

jimthompson5910 (jim_thompson5910):

ok, that's good it's not completely new as for the trig stuff, all of it seems completely foreign?

jimthompson5910 (jim_thompson5910):

hopefully you've seen some of the identities before

OpenStudy (anonymous):

all of the trig is new... and sadly i have another quiz and a test after the quiz i am working on... and i have to get 80 or above on all of them...

OpenStudy (anonymous):

@jim_thompson5910 you still there?

jimthompson5910 (jim_thompson5910):

yeah sorry I got distracted and I just noticed the notification as a rule we can't help with quizzes, tests, or exams of any kind (since they must be done on your own). But you can go to the hw, lectures or study guides to work on similar problems. We can help with those. Working on them will help you practice and get ready for the tests.

jimthompson5910 (jim_thompson5910):

When did they introduce all of this new trig? At the last minute so to speak?

OpenStudy (anonymous):

we are in unit 6 and this short unit is entirely trig so yes it was very last minute

jimthompson5910 (jim_thompson5910):

unit 6 is the last unit? I guess that would make sense (in a way) because they're trying to transition you from algebra to trig

jimthompson5910 (jim_thompson5910):

the transition is a bit too fast though

jimthompson5910 (jim_thompson5910):

how much do you have in terms of hw? or lecture material?

OpenStudy (anonymous):

it isnt the last unit. we have a couple units after this and, according to the names of the lessons, we are going back to things that i learned in algebra 1. so this unit is completely random

jimthompson5910 (jim_thompson5910):

for this particular unit

jimthompson5910 (jim_thompson5910):

hmm weird

jimthompson5910 (jim_thompson5910):

it's like they just gave up trying and clobbered together a whole bunch of things

OpenStudy (anonymous):

it is making it impossible for me to get my grade up enough so that i can pass and graduate (i am in my senior year of high school and i dont understand any of this)

jimthompson5910 (jim_thompson5910):

which is why I was hoping transferring was possible (so you can get into a class that actually teaches trig if it's a trig class...not a weird mix)

jimthompson5910 (jim_thompson5910):

anyways, I might as well finish up since I got this far so we have cos(theta) = -1/2 we use the unit circle http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/720px-Unit_circle_angles_color.svg.png to find values of theta that make cos(theta) = -1/2 true. Basically we are looking for points on the unit circle that have x coordinates of -1/2 According to the unit circle, cos(120) = -1/2 and cos(240) = -1/2 So theta = 120 and theta = 240 those angles are in degree mode The unit circle provided also shows radians as well. Because 120 degrees = 2pi/3 radians and 240 degrees = 4pi/3 radians, this means the solutions to the original equation are 2pi/3 & 4pi/3 There are infinitely many solutions, but remember theta is restricted to be in the interval from 0 to 2pi